# for Large Language Models

### v1

@@ Abstract
The capacity of Large Language Models (LLMs) to comprehend and reason over long contexts is pivotal for advancements in diverse fields. Yet, they still struggle with identifying relevant contexts and memory searching. To address this issue, we introduce Query-aware Inference for LLMs (Q-LLM), a system designed to process extensive sequences akin to human cognition. By focusing on memory data relevant to a given query, Q-LLM accurately captures pertinent information within a fixed window size and provides precise answers to queries. It requires no additional training and can be seamlessly integrated with any LLMs. Using LLaMA3 (QuickLLaMA), Q-LLM can read Harry Potter within 30 seconds and accurately answer related questions. On widely recognized benchmarks, Q-LLM improved performance by 7.17% compared to the current state-of-the-art on LLaMA3 and by 3.26% on Mistral on the $\infty$-bench. In the Needle-in-a-Haystack task, Q-LLM improved upon the current SOTA by 7.0% on Mistral and achieved 100% accuracy on LLaMA3. Our code can be found in https://github.com/dvlab-research/Q-LLM.
@@ 1 Introduction
Figure 1: Examples of our QuickLLaMA-8B (1) reading long context containing 100K tokens, (2) reading our paper that has not be seen in the pretrained dataset, (3) retrieving value in long key-value pairs and (4) retrieving in Needle-in-a-Haystack task. More examples and comparisons with the SOTAs are provided in Appendix E.
Figure 2: This is an example from the $\infty$-Bench. Three questions were posed about the same long book: (1) Which among Annalisa, Seb, Peyton, and Gannonmarie is not Mrs. Bronwyn’s child? (2) What’s the name of the Bronwyns’ summer home? (3) Who among Mrs. Bronwyn, Mrs. Deandra, Rosemarie, and Cael is the final to perish? We present the score heatmap of the first 50 memory blocks. The methods used include (a) the consistent results from InfLLM for all three queries, and (b-d) the query-aware results from Q-LLM.
@@
To address these challenges, we design Query-aware Inference for LLMs (Q-LLM), which processes extensive sequences in a manner similar to human cognition. Humans, when interpreting text, initially examine the question, and then seek the answer within the context, keeping the query in mind. This idea forms the foundation of our Query-aware Context Lookup strategy. Only memory data pertinent to the query is chosen for each computational step, disregarding unrelated distractions. As a result, LLMs can capture pertinent information within a fixed window size and provide precise answers to queries. Q-LLM doesn’t require extra training and can be seamlessly integrated with any LLMs.
We assess the performance of Q-LLM by utilizing LLaMA3-8B-inst AI@Meta (2024) (QuickLLaMA) and Mistral-7B-inst-v0.2 Jiang et al. (2023) as foundational models. These base models are pre-trained on sequences that do not exceed $8$K tokens. Instead, our QuickLLaMA can read Harry Potter with 100K tokens within half a minute on a single 800 GPU and accurately answer the questions, as shown in Fig. 1. We employ several widely recognized benchmarks, namely Longbench Bai et al. (2023), $\infty$-Bench Zhang et al. (2024), and Needle-in-a-Haystack Benchmark. Specifically, with a context window of 512, Q-LLM improved by 7.17% compared to the current SOTA on LLaMA3, and by 3.26% on Mistral on the $\infty$-bench. In the Needle-in-a-Haystack task, Q-LLM improved upon the current SOTA by by 7.0% on Mistral and achieves 100% on LLaMA3. We have extended the input sequence to contexts of 1024K length, further demonstrating our model’s capability in handling extremely long sequences.
@@ 2 Related Works
The issue of how LLMs process long sequences has been a significant topic of study for a long time and has been extensively researched Dong et al. (2023); Tay et al. (2022); Huang et al. (2023); Chen et al. (2023b). The approaches can generally be categorized into two main methods, our primary focus is on efficient context computation, which seeks to improve the computational efficiency of attention layers. This allows LLMs to make better use of limited resources when processing longer sequences. To ensure the completeness of the survey, in Appendix D, we provide an introduction to the other method, context length extrapolation. It aims to enable LLMs trained on short sequences to handle much longer sequences. We also mention the relevant work on memory-based models.
@@
The computational and memory demands of LLM training often limit it to short sequences. Using LLMs directly on long sequences presents challenges such as out-of-domain issues and distractions from lengthy and noisy inputs Lin et al. (2024); Tworkowski et al. (2024a); Li et al. (2024d). As a result, context length extrapolation has emerged as a method to extend LLMs’ sequence length without additional training. Early approaches have designed new relative positional encoding mechanisms during pre-training Press et al. (2022); Tworkowski et al. (2024b). The following research has focused on the extensively adopted rotary position embedding (RoPE) Su et al. (2023), suggesting extending the length by interpolating positions to introduce non-integer positions Chen et al. (2024); Peng et al. (2023); Jin et al. (2024); Chen et al. (2023a). While these approaches can mitigate the out-of-domain issue stemming from unseen lengths, they fall short in addressing the distracting challenge posed by noisy context. In response to this limitation, Stream-LLM (Xiao et al., 2024b) and LM-Infinite(Lin et al., 2024) utilize the sliding window attention mechanism, discarding distant contexts to seamlessly process extremely long sequences. Additionally, InfLLM Xiao et al. (2024a) leverages a context memory to furnish LLMs with pertinent contextual information. These techniques, however, offer limited enhancement in capturing long-distance dependencies crucial for understanding lengthy texts. The model’s objective during long-text reading is inherently ambiguous, and it can become distracting when reading extensive articles. In this work, we introduce the Query-aware Context Lookup mechanism, enabling the model to effectively retrieve information relevant to the query from lengthy texts.
Figure 3: The illustration of our Q-LLM framework. The input from the memory context is partitioned into memory blocks, which are searched by Query-aware Context Lookup for query-related blocks. The current key-value cache comprises global tokens, query tokens, query-related blocks, and local tokens. Together, these form a new context window that, along with current tokens, is fed into the LLM.
@@ 3 Methods
In this section, we introduce the overall framework of Query-aware Inference for LLMs (Q-LLM) in Sec. 3.1, as depicted in Fig. 3. Then, we propose the preliminary memory block in Sec. 3.2 and our proposed Query-aware Context Lookup in Sec. 3.3.
@@ 3.1 Overall Framework
An example of these tokens in the prompt is shown in Fig. 4. We propose the Query-aware Context Lookup strategy to find the query-related tokens $\mathbf{R}$ from the context tokens $\mathbf{C}$:
@@
where $\phi(\cdot)$ refers to the lookup operation of context memory. We will detail the strategy in Sec. 3.3. For each step, Q-LLM combines the global tokens, query tokens, query-related tokens, and local tokens to compose the current key-value cache.
@@
Method Infinite Stream Infllm Q-LLM Infinite Stream Infllm Q-LLM Infinite Stream Infllm Q-LLM
@@
Method Infinite Stream Infllm Q-LLM Infinite Stream Infllm Q-LLM Infinite Stream Infllm Q-LLM
@@
Table 1: The results comparison based on Mistral-7B-inst-v0.2 Jiang et al. (2023). The $95$% quantile for text lengths in these two benchmarks is $214$K and $31$K. Our results are highlighted in teal and best results are indicated in bold.
Method Infinite Stream Infllm Q-LLM Infinite Stream Infllm Q-LLM Infinite Stream Infllm Q-LLM
Context Window 512 1K 2K
En.MC 37.12 34.93 37.77 40.17 37.12 34.93 37.99 40.17 36.24 37.12 33.19 34.50
Retrieve.Number 3.39 3.39 96.61 98.98 3.39 3.39 40.68 41.19 3.39 3.39 28.64 27.63
Code.Debug 22.59 22.59 22.59 22.59 22.59 22.59 23.10 23.86 24.11 22.84 22.59 23.10
Math.Find 20.86 19.71 29.23 30.70 20.86 19.71 32.29 31.14 18.00 16.86 26.86 27.37
Retrieve.KV 0.80 0.80 24.40 61.20 0.80 0.80 57.20 70.80 1.80 1.80 80.80 84.00
Average 14.69 14.14 51.77 58.94 14.69 14.14 48.54 51.19 14.49 14.23 48.68 49.43
@@
Method Infinite Stream Infllm Q-LLM Infinite Stream Infllm Q-LLM Infinite Stream Infllm Q-LLM
Context Window 512 1K 2K
NarrativeQA 14.50 14.56 19.28 19.29 14.50 14.56 19.90 20.50 16.47 15.12 19.41 25.60
Qasper 21.06 20.72 26.08 26.58 21.06 20.72 32.35 31.47 32.01 31.72 41.27 39.12
MultiFieldQA 25.66 25.79 36.01 40.12 25.66 25.79 41.46 46.44 31.63 30.99 45.89 48.30
HotpotQA 31.95 32.84 41.42 42.34 31.95 32.84 43.75 49.15 34.73 35.26 44.97 49.91
2WikiMQA 24.72 24.28 28.44 29.63 24.72 24.28 30.83 31.53 29.22 30.59 36.27 39.63
Musique 12.72 13.62 17.48 18.75 12.72 13.62 21.26 23.95 13.50 13.64 19.73 25.03
GovReport 26.25 25.93 29.26 26.83 26.25 25.93 30.44 28.73 27.84 27.83 30.68 29.80
QMSum 19.38 19.42 19.10 19.02 19.38 19.42 20.30 20.62 19.91 20.14 21.36 22.23
MultiNews 26.42 26.57 26.63 25.22 26.42 26.57 27.46 26.85 27.36 27.37 27.87 27.85
TriviaQA 82.46 82.47 80.81 77.65 82.46 82.47 85.11 85.04 88.07 87.35 88.03 87.70
SAMSum 38.28 37.91 38.73 38.83 38.28 37.91 39.59 40.40 36.93 35.97 34.86 34.97
PassageRetrieval 13.75 13.25 19.50 35.00 13.75 13.25 58.75 69.00 23.50 23.50 85.25 88.00
LCC 54.17 53.88 53.24 52.39 56.32 53.88 60.23 60.43 60.42 58.15 58.17 58.37
RepoBench-P 60.71 60.59 58.33 57.04 62.81 60.59 60.86 60.51 64.95 62.97 62.01 61.04
Average 32.29 32.27 35.31 36.34 32.59 32.27 40.88 42.47 36.18 35.76 43.98 45.54
@@
Table 2: The results comparison based on LLaMA3-8B-inst AI@Meta (2024). The $95$% quantile for text lengths in these two benchmarks is $214$K and $31$K. Our results are highlighted in teal and best results are indicated in bold.
(a) Mistral-7B-inst-v0.2
(b) LLaMA3-8B-inst
Figure 5: The comparison of performance in the Needle-in-a-Haystack task. The horizontal axis represents the document’s length (the haystack), whereas the vertical axis specifies the location of a brief sentence (the needle) within the document, ranging from 1K to 128K tokens. For instance, a 50% value implies that the needle is situated halfway through the document. A red cell indicates the language model’s inability to recall the needle’s information, while a green cell denotes successful recall by the model.
@@ 4 Experiments
In this section, we conduct experiments utilizing Mistral-7B-inst-v0.2 Jiang et al. (2023) and LLaMA3-8B-inst AI@Meta (2024) as our base models. We compare our methods with competing sliding window approaches, containing LM-Infinite (Infinite) Lin et al. (2024), StreamingLLM (Stream) Xiao et al. (2024b) and InfLLM (infllm) Xiao et al. (2024a). We test the methods on three cache lengths: 512, 1024 (1K), and 2048 (2K). More configuration details are in Appendix C. We use the same prompts for all methods to ensure a fair comparison, as detailed in Appendix F.
Figure 6: Ablation of query weight $\beta$.
@@ 4.1 Long-Bench and $\infty$-Bench
In this section, we utilize representative tasks from two widely-recognized long document benchmarks, $\infty$-Bench Zhang et al. (2024) and Long-Bench Bai et al. (2023). These benchmarks encompass a variety of tasks including question answering, summarization, few-shot learning, context retrieval, mathematical computing, code completion, etc. The $95$% quantile for sequence lengths in $\infty$-Bench and Long-Bench reaches $214$K and $31$K tokens. The outcomes based on Mistral-7B-inst-v0.2 and LLaMA3-8B-inst are detailed in Tab. 1 and Tab. 2 respectively. The following observations can be made from the results: (1) Our approach shows considerable enhancement in performance when compared to models utilizing the sliding window mechanism (StreamingLLM and LM-Infinite) across benchmarks and context window lengths. This suggests that the context memory in Q-LLM can effectively provide LLMs with appropriate contextual data, facilitating efficient comprehension and reasoning on long sequences. (2) Our method also exhibits a significant performance uplift when compared to models with other lookup mechanisms (InfLLM). This implies that previous methods still struggle to extract valid information from noisy contexts. Our proposed Query-aware Context Lookup, however, can purposefully use the query to find relevant information in the long context. (3) Our technique is particularly beneficial in scenarios with longer input contexts and relatively smaller available context windows, as observed in comparisons across different benchmarks and context budgets. For instance, with a context window of 512 on the $\infty$-bench, Q-LLM improved by 7.17% compared to the current SOTA on LLaMA3, and by 3.26% on Mistral. This illustrates our model’s superiority in infinite stream scenarios. (4) Our model’s relative improvement is more noticeable on LLaMA3 when compared to other models. This is because superior models can more effectively utilize query information to precisely locate relevant information in the long context. This paves the way for our model to further enhance its inference capabilities with the development of superior LLMs in future.

### v2

@@ Abstract
The capacity of Large Language Models (LLMs) to comprehend and reason over long contexts is pivotal for advancements in diverse fields. Yet, they still struggle with identifying relevant contexts and memory searching. To address this issue, we introduce Query-aware Inference for LLMs (QLLM), a system designed to process extensive sequences akin to human cognition. By focusing on memory data relevant to a given query, QLLM accurately captures pertinent information within a fixed window size and provides precise answers to queries. It requires no additional training and can be seamlessly integrated with any LLMs. Using LLaMA3 (QuickLLaMA), QLLM can read Harry Potter within 30 seconds and accurately answer related questions. On widely recognized benchmarks, QLLM improved performance by 7.17% compared to the current state-of-the-art on LLaMA3 and by 3.26% on Mistral on the $\infty$-bench. In the Needle-in-a-Haystack and BABILong task, QLLM improved upon the current SOTA by 7.0% and 6.1%. Our code can be found in https://github.com/dvlab-research/Q-LLM
@@ 1 Introduction
Figure 1: Examples of our QuickLLaMA-8B (1) reading long context containing 100K tokens, (2) reading our paper that has not be seen in the pretrained dataset, (3) retrieving value in long key-value pairs and (4) retrieving in Needle-in-a-Haystack task. More examples and comparisons with the SOTAs are provided in Appendix D.
Figure 2: This is an example from the $\infty$-Bench. Three questions were posed about the same long book: (1) Which among Annalisa, Seb, Peyton, and Gannonmarie is not Mrs. Bronwyn’s child? (2) What’s the name of the Bronwyns’ summer home? (3) Who among Mrs. Bronwyn, Mrs. Deandra, Rosemarie, and Cael is the final to perish? We present the score heatmap of the first 50 memory blocks. The methods used include (a) the consistent results from InfLLM for all three queries, and (b-d) the query-aware results from QLLM.
@@
To address these challenges, we design Query-aware Inference for LLMs (QLLM), which processes extensive sequences in a manner similar to human cognition. Humans, when interpreting text, initially examine the question, and then seek the answer within the context, keeping the query in mind. This idea forms the foundation of our Query-aware Context Lookup strategy. Only memory data pertinent to the query is chosen for each computational step, disregarding unrelated distractions. As a result, LLMs can capture pertinent information within a fixed window size and provide precise answers to queries. QLLM doesn’t require extra training and can be seamlessly integrated with any LLMs.
We assess the performance of QLLM by utilizing LLaMA3-8B-inst AI@Meta (2024) (QuickLLaMA) and Mistral-7B-inst-v0.2 Jiang et al. (2023) as foundational models. These base models are pre-trained on sequences that do not exceed $8$K tokens. Instead, our QuickLLaMA can read Harry Potter with 100K tokens within half a minute on a single A800 GPU and accurately answer the questions, as shown in Fig. 7. We employ several widely recognized benchmarks, namely Longbench Bai et al. (2023), $\infty$-Bench Zhang et al. (2024), and Needle-in-a-Haystack Benchmark. Specifically, with a context window of 512, QLLM improved by 7.17% compared to the current SOTA on LLaMA3, and by 3.26% on Mistral on the $\infty$-bench. In the Needle-in-a-Haystack task, QLLM improved upon the current SOTA by by 7.0% on Mistral and achieves 100% on LLaMA3. In the BABILong task, QLLM improved upon the current SOTA by 6.1%. We have extended the input sequence to contexts of 1048K length, further demonstrating our model’s capability in handling extremely long sequences.
Figure 3: The illustration of our QLLM framework. The input from the memory context is partitioned into memory blocks, which are searched by Query-aware Context Lookup for query-related blocks. The current key-value cache comprises global tokens, query tokens, query-related blocks, and local tokens. Together, these form a new context window that, along with current tokens, is fed into the LLM.
@@ 2 Related Works
Efficient Context Computation. The computational and memory demands of LLM training often limit it to short sequences. Using LLMs directly on long sequences presents challenges such as out-of-domain issues and distractions from lengthy and noisy inputs Lin et al. (2024); Tworkowski et al. (2024a); Li et al. (2024d). As a result, context length extrapolation has emerged as a method to extend LLMs’ sequence length without additional training. Early approaches have designed new relative positional encoding mechanisms during pre-training Press et al. (2022); Tworkowski et al. (2024b). The following research has focused on the extensively adopted rotary position embedding (RoPE) Su et al. (2023), suggesting extending the length by interpolating positions to introduce non-integer positions Chen et al. (2024); Peng et al. (2023); Jin et al. (2024); Chen et al. (2023). To process extremely long sequences, Stream-LLM (Xiao et al., 2024b) and LM-Infinite(Lin et al., 2024) utilize the sliding window attention mechanism and discard distant contexts. Additionally, InfLLM Xiao et al. (2024a) leverages a context memory to furnish LLMs with pertinent contextual information. Yet, the objective of these models during long-text reading is inherently ambiguous, and it can become distracting when reading extensive articles. In this work, we introduce the Query-aware Context Lookup mechanism, enabling the model to effectively retrieve information relevant to the query from lengthy texts.
Context Length Extrapolation. The computational complexity of attention layers, which grows quadratically, is a significant bottleneck restricting LLMs’ capability to handle lengthy sequences. Consequently, numerous researchers have devised efficient attention mechanisms, including sparse attention Zaheer et al. (2021); Beltagy et al. (2020); Child et al. (2019); Ainslie et al. (2020); Zhao et al. (2019), approximate attention computations using kernel functions Kitaev et al. (2020); Wang et al. (2020); Katharopoulos et al. (2020), and replacing attention layers with state-space models of linear complexity Gu et al. (2022); Gu and Dao (2023). These approaches necessitate architectural modifications, requiring retraining of the models. Concurrently, many scholars have tackled this challenge from an infrastructural angle by optimizing the memory usage of attention computations to mitigate the computational resource requirements of the model Dao et al. (2022); Dao (2023); Hong et al. (2024); Shazeer (2019); Kwon et al. (2023). Given the training-free nature of our method, it can be seamlessly integrated to further expedite LLM inference.
Memory-based Approaches. Memory networks have been extensively researched for decades and have demonstrated effectiveness in enhancing models with additional information storage capabilities Graves et al. (2014); Weston et al. (2015); Sukhbaatar et al. (2015); Miller et al. (2016). With the rise of pre-trained models, memory layers have gradually found application in the training stage of recurrent transformer layers, enabling models to recursively process long sequences Dai et al. (2019); Rae et al. (2020); Khandelwal et al. (2020); Wu et al. (2022); Bertsch et al. (2023). These approaches segment sequences, encoding each segments individually, and utilize memory to retain context information from preceding segments. Yet, they necessitate architectural modifications and are typically incorporated during the pre-training phase. In contrast, our objective is to explore the intrinsic properties of LLMs and introduce a training-free Query-aware Context Lookup mechanism for long-text comprehension.
@@ 3 Methods
In this section, we introduce the overall framework of Query-aware Inference for LLMs (QLLM) in Sec. 3.1, as depicted in Fig. 3. Then, we propose the preliminary memory block in Sec. 3.2 and our proposed Query-aware Context Lookup in Sec. 3.3.
@@ 3.1 Overall Framework
An example of these tokens in the prompt is shown in Fig. 4. Given that all memory blocks are necessary to be maintained and most of them and seldom used, we adopt an offloading strategy, which stores most memory blocks in CPU memory. More details are in Sec. A.3. We propose the Query-aware Context Lookup strategy to find the query-related tokens $\mathbf{R}$ from the context tokens $\mathbf{C}$:
@@
where $\phi(\cdot)$ refers to the lookup operation of context memory. We will detail the strategy in Sec. 3.3. For each step, QLLM combines the global tokens, query tokens, query-related tokens, and local tokens to compose the current key-value cache.
@@
Method Infinite Stream ILM QLLM Infinite Stream ILM QLLM Infinite Stream ILM QLLM
@@
Method Infinite Stream ILM QLLM Infinite Stream ILM QLLM Infinite Stream ILM QLLM
@@
Table 1: The results comparison based on Mistral-7B-inst-v0.2 Jiang et al. (2023). Our results are highlighted in teal and best results are indicated in bold.
Method LLaMA Infinite Stream ILM QLLM Infinite Stream ILM QLLM Infinite Stream ILM QLLM
Context Window -1048K 512 1K 2K
En.MC 31.0 37.12 34.93 37.77 40.17 37.12 34.93 37.99 40.17 36.24 37.12 33.19 34.50
Retrieve.PassKey 6.78 3.40 3.40 100.0 100.0 3.40 3.40 100.0 100.0 3.40 3.40 100.0 100.0
Retrieve.Number 6.78 3.39 3.39 96.61 98.98 3.39 3.39 40.68 41.19 3.39 3.39 28.64 27.63
Code.Debug 22.59 22.59 22.59 22.59 22.59 22.59 22.59 23.10 23.86 24.11 22.84 22.59 23.10
Math.Find 34.29 20.86 19.71 29.23 30.70 20.86 19.71 32.29 31.14 18.00 16.86 26.86 27.37
Retrieve.KV 6.2 0.80 0.80 24.40 61.20 0.80 0.80 57.20 70.80 1.80 1.80 80.80 84.00
Average 17.94 14.69 14.14 51.77 58.94 14.69 14.14 48.54 51.19 14.49 14.23 48.68 49.43
@@
Method LLaMA Infinite Stream ILM QLLM Infinite Stream ILM QLLM Infinite Stream ILM QLLM
Context Window -1048K 512 1K 2K
NarrativeQA 23.78 14.50 14.56 19.28 19.29 14.50 14.56 19.90 20.50 16.47 15.12 19.41 25.60
Qasper 21.22 21.06 20.72 26.08 26.58 21.06 20.72 32.35 31.47 32.01 31.72 41.27 39.12
MultiFieldQA 39.89 25.66 25.79 36.01 40.12 25.66 25.79 41.46 46.44 31.63 30.99 45.89 48.30
HotpotQA 17.16 31.95 32.84 41.42 42.34 31.95 32.84 43.75 49.15 34.73 35.26 44.97 49.91
2WikiMQA 18.11 24.72 24.28 28.44 29.63 24.72 24.28 30.83 31.53 29.22 30.59 36.27 39.63
Musique 10.39 12.72 13.62 17.48 18.75 12.72 13.62 21.26 23.95 13.50 13.64 19.73 25.03
GovReport 33.76 26.25 25.93 29.26 26.83 26.25 25.93 30.44 28.73 27.84 27.83 30.68 29.80
QMSum 23.38 19.38 19.42 19.10 19.02 19.38 19.42 20.30 20.62 19.91 20.14 21.36 22.23
MultiNews 27.68 26.42 26.57 26.63 25.22 26.42 26.57 27.46 26.85 27.36 27.37 27.87 27.85
TriviaQA 87.76 82.46 82.47 80.81 77.65 82.46 82.47 85.11 85.04 88.07 87.35 88.03 87.70
SAMSum 41.89 38.28 37.91 38.73 38.83 38.28 37.91 39.59 40.40 36.93 35.97 34.86 34.97
PassageRetrieval 51.83 13.75 13.25 19.50 35.00 13.75 13.25 58.75 69.00 23.50 23.50 85.25 88.00
LCC 43.79 54.17 53.88 53.24 52.39 56.32 53.88 60.23 60.43 60.42 58.15 58.17 58.37
RepoBench-P 46.11 60.71 60.59 58.33 57.04 62.81 60.59 60.86 60.51 64.95 62.97 62.01 61.04
Average 34.77 32.29 32.27 35.31 36.34 32.59 32.27 40.88 42.47 36.18 35.76 43.98 45.54
@@
Table 2: The results comparison based on LLaMA3-8B-inst AI@Meta (2024). Our results are highlighted in teal and best results are indicated in bold.
@@ 3.3 Query-aware Context Lookup
Figure 5: The comparison of performance in the Needle-in-a-Haystack task. The horizontal axis represents the document’s length (the haystack), whereas the vertical axis specifies the location of a brief sentence (the needle) within the document, ranging from 1K to 128K tokens. A red cell indicates the language model’s inability to recall the needle’s information, while a green cell denotes successful recall by the model.
@@ 4 Experiments
In this section, we conduct experiments utilizing Mistral-7B-inst-v0.2 Jiang et al. (2023) and LLaMA3-8B-inst AI@Meta (2024) as our base models. We compare our methods with LLaMA3-8B-inst-1048K (LLaMA-1048K) Gradient (2024) and other competing sliding window approaches, containing LM-Infinite (Infinite) Lin et al. (2024), StreamingLLM (Stream) Xiao et al. (2024b) and InfLLM (Infllm) Xiao et al. (2024a). We test the methods on three cache lengths: 512, 1024 (1K), and 2048 (2K). More configuration details are in Appendix C. Note that we add the queries before the context to ensure the baselines also have query-aware capabilities, as detailed in Appendix E.
Figure 6: (a) Memory and (b) Time consumption of different methods as tokens increases. (c) Ablation of query weight $\beta$. (d) The results of methods on sequences with extremely lengthy sequence lengths.
@@ 4.1 Long-Bench and $\infty$-Bench
In this section, we utilize representative tasks from two widely-recognized long document benchmarks, $\infty$-Bench Zhang et al. (2024) and Long-Bench Bai et al. (2023). The $95$% quantile for sequence lengths in $\infty$-Bench and Long-Bench reaches $214$K and $31$K tokens. The outcomes based on Mistral-7B-inst-v0.2 and LLaMA3-8B-inst are detailed in Tab. 1 and Tab. 2 respectively. The following observations can be made from the results: (1) Our approach shows considerable enhancement in performance when compared to base models (LLaMA3-8B-inst-1048K ) and that utilizing the sliding window mechanism (StreamingLLM and LM-Infinite) across benchmarks and context window lengths. This suggests that the context memory in QLLM can effectively provide LLMs with appropriate contextual data, facilitating efficient comprehension and reasoning on long sequences. (2) Our method also exhibits a significant performance uplift when compared to models with other lookup mechanisms (InfLLM). This implies that previous methods still struggle to extract valid information from noisy contexts. Our proposed Query-aware Context Lookup, however, can purposefully use the query to find relevant information in the long context. (3) Our technique is particularly beneficial in scenarios with longer input contexts and relatively smaller available context windows, as observed in comparisons across different benchmarks and context budgets. For instance, with a context window of 512 on the $\infty$-bench, QLLM improved by 7.17% compared to the current SOTA on LLaMA3, and by 3.26% on Mistral. This illustrates our model’s superiority in infinite stream scenarios. (4) Our model’s relative improvement is more noticeable on LLaMA3 when compared to other models. This is because superior models can more effectively utilize query information to precisely locate relevant information in the long context.